Determine the area of the normal distribution with mean of 10, standard deviation of 5 and scores between 5 to 12



1
Expert's answer
2022-04-10T16:32:47-0400

We have a normal distribution, "\mu=10, \sigma=5."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"z_1=\cfrac{5-10}{5}=-1;\\
z_2=\cfrac{12-10}{5}=0.4;\\
P(5<X<12)=P(-1<Z<0.4)=\\
=P(Z<0.4)-P(Z<-1)=\\
=0.6554-0.1587=0.4967 \text{ (from z-table).}"